OpenStudy (anonymous):

The graph of a line passes through the two points (-2, 1) and (2, 1). What is the equation of the line written in general form? x - y + 1 = 0?

OpenStudy (anonymous):

do you know the slope of the line that passes through the 2 points?

OpenStudy (anonymous):

no graph was given...so no i dont :/

OpenStudy (anonymous):

... IMPORTANT LINE RELATED EQUATIONS TO KNOW AND MEMORIZE slope formula m= slope/ gradiant -- same thing \[m=\frac{y_2-y_1}{x_2-x_1}\] standard formula \[Ax+By=C\] point-slope formula \[y-y_1=m(x-x_1)\] slope-intercept formula b= y-intercept -- in the form of (0,y) \[y=mx+b\]

OpenStudy (anonymous):

now what's the slope?

OpenStudy (anonymous):

y=2/1?

OpenStudy (anonymous):

you're solving for m....

OpenStudy (anonymous):

oh :/

OpenStudy (anonymous):

m is slope... ALWAYSSSS

OpenStudy (anonymous):

hmm....

OpenStudy (anonymous):

Y-1=0?

OpenStudy (anonymous):

good enough y=1

OpenStudy (anonymous):

Yes!! lol

OpenStudy (anonymous):

honestly just memorize those 4 simple equations and your life will be so much easier WHICH MEANS YOU SHOULD READ THROUGH ALL OF IT

OpenStudy (anonymous):

Ok, I'll try (: Thank You!